OpenStudy (christos):

Express 3702_8 in decimal

OpenStudy (unklerhaukus):

\[3702_8\\ =3\times8^3+7\times8^2+0\times8^1+2\times8^0\\ =3\times512+7\times64+2\times1\\ =1536+448+2\\ =1986\]